SPIN-SPIN COUPLING CONSTANTS ACROSS N-H^{+}-N HYDROGEN BONDS

\underline{J. E. DEL BENE}, Department of Chemistry, Youngstown State University, Youngstown, Ohio 44555; J. ELGUERO, Instituto de Química Médica,CSIC, Juan de la Cierva, 3, E-28006 Madrid, Spain.

A systematic ab initio EOM-CCSD study of ^{15}N-^{15}N and ^{15}N-^{1}H spin-spin coupling constants has been carried out for a series of complexes formed from 11 nitrogen bases with experimentally measured proton affinities. When these complexes are arranged in order of increasing proton affinity of the proton-acceptor base, and for each proton acceptor, increasing order of proton affinity of the protonated N-H donor, trends in distances and signs of coupling constants are evident which are indicative of the nature of the hydrogen bond. All two-bond spin-spin coupling constants (^{2h}J_{N−N}) are positive, and decrease as the N-N distance increases. All one-bond N-H coupling constants (^{1}J_{N−H}) are negative (^{1}K_{N−H} are positive). 1J_{N−H} is related to the N-H distance and the hybridization of the donor N atom. One bond H...N coupling constants (^{1h}J_{H−N}) are positive (^{1h}K_{H−N} are negative) for traditional hydrogen bonds, but ^{1h}J_{H−N} becomes negative when the hydrogen bond acquires sufficient proton-shared character. The N-N and H...N distances at which ^{1h}J_{H−N} changes sign are approximately 2.71 and 1.62 Å, respectively. Predictions are made of the values of ^{2h}J_{N−N} and ^{1}J_{N−H}, and the signs of ^{1h}J_{H−N} for those complexes that are too large for EOM-CCSD calculations.